OpenStudy (314dokg):

help me find the "t".. im sorry im quite confuse. d = vt + gt^2/2

OpenStudy (314dokg):

i'm kinda stuck at this part. 2d/2vgt = t

OpenStudy (owlcoffee):

Does it look like this?: \[d=\frac{ vt+g t ^{2} }{ 2 }\]

OpenStudy (anonymous):

Does this problem state that initial velocity is 0?

OpenStudy (314dokg):

nope... \[d = vt + \frac{ g t ^{2} }{ 2 }\] the velocity is 12.4 m/s... and this problem is more on projectile motion

OpenStudy (anonymous):

Can you copy and paste the whole problem?

OpenStudy (anonymous):

it will have 2 solutions for t second degree equation.

OpenStudy (314dokg):

1.) A dart player throws a dart horizontally at a speed of 12m/s. The dart hits the board 0.32 m below the height from which it was thrown -a.) How long did the dart travel? -b.) How far away is the player from the board? so far i wrote down the given: Vox = 12.4m/s ; dy = -0.32m i'm not even sure with the given

OpenStudy (anonymous):

put it in the form \[a x ^{2}+b x+c = 0\] and use the formula \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (314dokg):

i already converted some. all i need is just to find the y

OpenStudy (314dokg):

quadratic formula?

OpenStudy (anonymous):

Hey man v0 is zero. 

OpenStudy (314dokg):

umm... the vy is 0... since it says horizontal... it means its traveling on the x axis...

OpenStudy (anonymous):